OpenStudy (anonymous):

What is the sum of the arithmetic sequence 3, 11, 19 ..., if there are 26 terms?

OpenStudy (anonymous):

common difference is 8

OpenStudy (anonymous):

first term is 3

OpenStudy (anonymous):

3x + 8x (x - 1)/2 x = the number of terms so 26.

OpenStudy (amistre64):

there are a few ways to approach this, either way you have to determine the a1 + d(n-1) setup the 26th term is of course: 3 + 8(25) 26 [3+3+8(25)] --------------- 2

OpenStudy (amistre64):

26 terms first = 3 last = 3+8(25) first + last = 3+3+8(25) = 206 26(206) = 5356 and half of that is ....

OpenStudy (anonymous):

2,678! Thank You! :)

OpenStudy (amistre64):

it gets easier with practice :) your welcome